# Efficient Monte Carlo Event Generation for Neutrino-Nucleus Exclusive Cross Sections

### v2

@@ V Discussion
The greater challenge arises in describing $2p2h$ processes due to the increased number of dimensions in the phase space. These processes correspond to two-nucleon knockout, and their fully exclusive differential cross-sections involve nine independent kinematic variables:
@@ VI Conclusion
Looking ahead, the flexibility of normalizing flows makes them well-suited to other processes beyond the $1p1h$ process studied here. In particular, higher-dimensional processes such as $2p2h$ or single-pion production can benefit from a similar surrogate approach. Currently, neutrino experiments like T2K rely on “Frankenstein” models that stitches together submodels based on different theoretical frameworks, each tuned to a specific process, and combined with ad-hoc relative proportions. In contrast, a normalizing flow-based surrogate can make possible a unified sampling scheme for all processes within a single, theory-grounded cross section model, such as the Mean Field cross section model developed by the Ghent group. Such a tool could be crucial for future neutrino experiments, delivering fast, accurate, and cohesive event generation that seamlessly bridges the gap between the experimental and theoretical frontiers.

### v3

@@ V Discussion
The greater challenge arises in describing $2p2h$ processes due to the increased number of dimensions in the phase space. These processes correspond to two-nucleon knockout, and their fully exclusive differential cross sections involve nine independent kinematic variables:
@@ VI Conclusion
Looking ahead, the flexibility of normalizing flows makes them well-suited to other processes beyond the $1p1h$ process studied here. In particular, higher-dimensional processes such as $2p2h$ or single-pion production can benefit from a similar surrogate approach. Currently, neutrino experiments such as T2K rely on so-called ”Frankenstein” models that are combinations of submodels drawn from different theoretical frameworks, each tuned to specific interaction channels and combined using ad-hoc relative proportions. While practical, this approach makes it difficult to disentangle whether observed features in predictions arise from genuine physical effects, differences in underlying frameworks, or purely phenomenological adjustments. In contrast, a normalizing flow-based surrogate can make a unified sampling scheme for all processes within a single, theory-grounded cross section model possible. It could do so even in the case of a mean-field model, such as that developed by the Ghent group, which has been before too computationally intensive to integrate directly into event generators in its full complexity. Such a tool could be crucial for future neutrino experiments, delivering fast, accurate, and cohesive event generation that seamlessly bridges the gap between the experimental and theoretical frontiers.
@@ Acknowledgments
Appendix A Profile of the proton kinetic energy for different angles of the outgoing muon and proton
To complement the analysis presented in Section IV.3.2, this appendix illustrates the impact of the biases between the true and surrogate cross sections on the distribution of proton kinetic energy $T_{N}$ at fixed kinematics, both with and without local averaging. The study is performed using a monoenergetic neutrino beam with energy $E_{\nu}=650\,\text{MeV}$, and the total cross section is computed by summing the contributions from the two occupied shells, as described in Equation 4. In the case of CCQE interaction in the impulse approximation, one can simply derive the kinetic energy by computing :
$T_{N}=\omega-E_{b}$
where $E_{b}$ is the shell dependent bounding energy.
Figures 6, 7 and 8 show the comparisons between true and surrogate cross sections for $\theta_{N}=5^{\circ}$, $20^{\circ}$, and $50^{\circ}$, respectively. In each figure, the top panel displays the differential cross section as a function of $T_{N}$ for various values of $\theta_{\mu}$, comparing the NF surrogate to the true differential cross section. The corresponding relative difference $\Delta$ is also shown. In the bottom panel, the same comparison is shown after averaging the cross section over a local 4D kinematic cube centered at the same point in $(T_{N},\theta_{\mu},\theta_{N},\phi_{N})$ space, using a bin size of $1$ MeV in $T_{N}$ and $1^{\circ}$ in each angular variable.
The top panels reveal small but noticeable differences between the true cross section and the prediction from the Normalizing Flows, both in the tail and in the core of the distribution. However, most of these differences, particularly in the core, are significantly reduced when averaging over a small neighborhood in kinematic space as illustrated in the bottom plots. This supports the conclusion that the remaining differences primarily reflect localized fluctuations rather than global modeling errors. Smaller deviations persist after averaging particularly in regions of low cross section. For analyses that probe regions of very low cross section (e.g., backward-going protons with respect to the momentum transfer or low $T_{N}$) or require sub-degree angular precision or sub-MeV energy resolution, the surrogate model may introduce non-negligible bias, making it necessary to reweight the events using the true cross section.
At the same time, the surrogate retains distinctive kinematic patterns predicted by the Ghent model, inherited from the shell modeling and the distorted wave approach. Using a generator whose truth distribution is matched to that surrogate model will therefore show these signatures in simulated samples, increasing the chances that experiments with near-$4\pi$ angular acceptance and high‑granularity detectors, with good sensitivity to low-momentum protons, such as the SuperFGD at T2K [56], will be able to observe some of them, if they are not already visible in existing datasets.
Figure 6: Profile of the proton kinetic energy distribution $T_{N}$ for fixed values of the muon scattering angle $\theta_{\mu}\in\{10^{\circ},20^{\circ},50^{\circ},80^{\circ},120^{\circ},170^{\circ}\}$, and for fixed outgoing nucleon angles $(\theta_{N},\phi_{N})=(5^{\circ},180^{\circ})$. The cross section is computed for a monoenergetic neutrino beam of $650$ MeV and for both shells combined. The top panel displays the differential cross section as a function of $T_{N}$ for each $\theta_{\mu}$ value, comparing the output of the NF surrogate to the true model. The NF prediction is shown with dashed lines, and the true model with solid lines. The relative difference $\Delta$ between the two curves is also shown. In the bottom panel, the same comparison is shown after averaging the cross section over a local 4D kinematic cube centered at the same $(T_{N},\theta_{\mu},\theta_{N},\phi_{N})$ point. The bin resolution of this cube is $1$ MeV in $T_{N}$ and $1^{\circ}$ in each angular dimension. Each cube contains $500$ uniformly distributed events sampled from the NF or from the true cross section.
Figure 7: $\theta_{N}=20^{\circ}$
Figure 8: $\theta_{N}=50^{\circ}$.